FEL Upgrade Project Weekly Brief – October 21-25, 2002


We have had another very busy and very successful week with installation, testing and commissioning activities for the FEL Upgrade project. (1) With successful installation and pump-out of the "x-chamber" and the shielded bellows assembly in the injector area, the entire linac is now under vacuum from the gun chamber to the end of the last cryomodule; (2) the "y-camber" at the end of the linac has been vacuum qualified and is ready for installation at the end of the linac next week along with its associated three dipole magnets (GU's,GV's); (3) the Personnel Safety System (PSS) was certified to allow SRF commissioning of the new cryomodule.; (4) the completed dipole magnets for the first recirculation arc reverse bend were inspected at the manufacturer to authorize shipment to us; and (5), the full electrode structure for the gun was dry assembled this week.


FEL Project management were presenters and participants in this week's DOE/SURA review of the Jefferson Lab's Institutional Management. The review was chaired by Dr. Charles Shank, Director of Lawrence Berkeley National Lab. From the preliminary assessment of the panel, the lab received good marks.

We continued with our final preparations for next week's "2002 Industrial Physics Forum", which Jefferson Lab is hosting for the American Institute of Physics. We are expecting 130 attendees who will hear about industrial applications of the FEL in a forum theme session entitled "New Tools for New Materials". The full program is given at www.aip.org/ipf.

Presentations were prepared for the Fall DEPS Symposium at Monterey, CA, on Nov. 12-14th.

Two talks were sent in to the Navy for review: an overview of the Jlab FEL program and a summary of FEL power scaling.

Next week we are looking forward to hosting a visit to JLab by Gil Graff and John Albertine from ONR on Thursday, Oct. 31 to discuss the FY02 cost and schedule performance and FY03 planning for the FEL project.

WBS 4 (Injector):
The shielded bellows connecting the gun chamber to the gate valve/lightbox was installed and the chamber pumped back down. The bellows/BPM was installed down stream off the lightbox and the region up to the cryounit was pumped down. A test fit up of all components in the ceramic stack was done minus the cesium channels. All components fit and works well. Cesium channels will be added into the ball cathode next week prior to installation onto the gun chamber.

WBS 6 (RF):
Zone 4 - This week the balance of the waveguides were installed, all leaks fixed, some wiring was corrected and the PSS Certification completed for this zone. The stepper motor connector for cavity 3 is broken and must be replaced. The 8 arc detectors for the cryomodule need to be re-installed.

Quarter HVPS - The software is nearly operational, but needs the status screens updated. No progress by made this week on unit 4, worked on zone 4.

Gun HVPS - The tank was pressurized to 14 psi. A leak was found in one of the HV drive connectors. The connector will be replaced next week.

Gun Resistor - The initial fit-up of the gun resistor mechanism was done this week. Several adjustments are being made by the Machine Shop and the assembly will continue next week.

WBS 8 (Instrumentation):
Drive Laser Controls - The new '8' channel (OMS VME58) stepper motor control card for the variable attenuator is installed and the modifications to the cabling are done as well. System check-out is in progress. The changes to the MPS shutter control chassis are done. The modifications were done to properly match the 24V relay logic to the Acromag I/O card. This system as well is being tested.

Beam viewers - The new PCB (F0100) for the beam viewer camera/lamp interface board is back from the board house. We were running short so took the opportunity to improve the design so assembly is easier. Beam viewers in the 0F through 2F region have been installed. All viewers that were to be installed this week have been tested. A few cables had to be repaired, and some of the beam viewer control boxes had to be repaired. The control/power cables for the same viewers were rang out and labeled yesterday. The wiggler viewers are also being assembled with the new control boxes and camera assemblies. Alignment will begin next week with the help from Z. Kursun from Ops.

The deep rack finally arrived for FL01B09 and was installed. The Optical klystron 5 kW PS was readily installed into the rack and the electricians are running AC power. A shunt chassis for magnet MGU5F10 has been installed in FL01B08. Magnet thermal interlock cables have been terminated. Work has started on the Ops screens for the Extraction Dipole PS. The software for trim and corrector magnets in the 1F and 2F regions are being developed and the screens updated with completion projected for late next week.

Modifications to the Dump Skid chassis for the new Solu Comp Conductivity Analyzer has been has been completed. Software for the BLM's (LeCroy HV mainframe) has been updated, compiled and ready for check out. Screens for the FSD have been updated to include the RF and P1 in Zone 4 and checked out. The optical cavity stepper motor chassis was sent to the fab four (EES assembly techs) for assembly. We are shooting to have the optical cavity external cables in place in one week.

We hosted a visit from Dr. Steve Smith of SLAC. Steve is an expert on Beam Position Monitoring systems. He advised us on techniques for the multipass BPM electronics that see both the accelerated and decelerated beams. We also had lengthy discussions on the 8 button BPMs as well as the regular strip lines.

DSP daughter board software: Presently we have a complete routine which has clk_in generation which involves generation of a pulse train of 500kHz (min off the board and the max the line scan pyro camera can handle; also the conversion times are better available), upon user request (like pressing of a push button on the board provided); detecting a successful rise or fall. Accordingly a pulse i.e. a high or opening of a GPIO switch is enabled. Upon successfully intercepting a high from the video signal, intertwined with the working of the clk_high or clk_low and the position of pulse_high, the ADC starts conversion and outputs the values to the console to be viewed. We have tested the same with single ended inputs and the error rates as well. Working upon integrating routines which determine the present state of the signals, presently this part of the code needs minimizing for better and proper tracking of variables.

WBS 9 (Transport):
Optical Chicane Dipoles (GW)
• The Purcell Gap sheet pieces are glued to 5 bottoms of the six magnets. The tops are at the machine shop for tooling ball hole machining.
• The comparison test stand is showing intermittent readings on the GWs showing the feeble nature of this legacy system. We are developing a plan to do all the readings on the stepper stand.
Injector Dipoles (DU/DV)
• The first GU was 2 % high in field integral with respect to the core field. We machined the pole face ends and the first magnet is being measured again. Field clamps will be adjusted next Monday for what we hope will be the qualifying measurement.
• The last GV in the injection string (for placement in the recirculating chicane) and the remaining two GVs for the Extraction Chicane have measurements which qualify them for FEL use and the papers have been sent to David Douglas for signing.
Arc 180 Degree Dipoles (GY)
• After a trip to Wang NMR they are now committed to working full time on starting potting the first coil and gluing Purcell Gap Shims within two weeks with final assembly and shipment of the first magnet projected for 4.5 weeks.
• We worked on some nagging problems. The coil can be squeezed to the print thickness but with pressures that give one pause for fear of developing shorts. We agreed to a larger thickness by .040 inches. This eats into the play room that we left for bedding the coil. However, their mold is so robust and flat (made of steel) that waviness of the coil will not be an issue. If the coil turns out to be too thick, it is mounted on spacer strips that can be made thinner.
• A second problem is the coil inner bend down radius is not (can’t ever be) the ideal radius per the drawings. The real bend is faired out either side of the center part of the radius, blending into the coil over an additional zone 3/4 long and 1/8 high. Forcing the coil to conform to the ideal radius will cause a short circuit in the coil. This blend is caused by the realities of the winding process and can’t be avoided. We will chamfer the mold edge by 1/8 inch x 3/4 inch where the conductor blends into the straight main body of the coil and similarly relieve the edge of the core’s edge at the coil zone.
Arc Bend, Reverse Bend Dipoles (GQ, GX)
• The 4th GX coil to complete the set of coils for the first arc was successfully impregnated.
• Inspection of the first two GQs and one GX revealed a few points that needed further work. They now plan to ship the four magnets in the middle of next week.
• Wang NMR plans to send the third GX and the third GQ along with the shipment of the first GY.
Trim Quad (QT)
• We now have three of the ten magnets left to measure.
Sextupole (SF)
• Master Machine is making the shim parts and will receive the original SCs next week for disassembly.
Octupole (OT)
• We have identified steel for manufacture and are working on a package.
Beam Line and Vacuum
• We are now on a production status where one arc chamber a week will come off the production line with a three week manufacturing time per chamber.
• Injection line is now under vacuum.
• Y Chamber passed tests and is about to be installed when the magnets are ready.
• Chicane chambers continue fit up, welding and machining
• The girder drawings for around the wiggler region were signed off
• We grouted the remaining stands
• The magnet measurement crew continued installation of the correctors, concentration on the injector region.

WBS 11 (Optics):
This has been a busy time as we prepare to install the optical cavity hardware in the vault. Since last week's report, the upper assembly of the optical cavity vacuum vessels (the top hats) were separated from the lower section while inside a HEPA filtered enclosure. The top hats will have parts welded to them, be leak-checked, and then be cleaned and ready for reassembly in about a week. The foundation plates that hold all of the optical cavity internals were cleaned for UHV, along with the rails and leadscrews. Assembly in the clean environment begins next week. The remaining cavity internals were sent out for cleaning and will be returned to us early next week. We have extended the clean area so we can build both of the internals and have one vacuum vessel at a time inside it. LVDTs are being prepped (cleaned and baked) for UHV use. A step forward in making them outgas at a low rate was made this week, when we learned how to remove a Viton spacer. At this writing, we appear to be on track to have the Survey and Alignment group align the cavity internals by the end of next week. The move downstairs would occur about one week after that.

Two new ultraviewers were assembled and are ready for integration on the support girder. The pedestals that hold the girder were grouted in place this week. The upstream OCMMS beamline hardware is installed, the second one will be installed after the entire optical klystron vacuum assembly is pumped down and leak-checked. Today, Survey and Alignment is marking the beam centerline on the floor near the optical cavity hardware positions to facilitate placement of our alignment lasers.

Progress continues in detailing and checking the prints for the optical transport turning mirrors (turning cassettes 1, 3, & 4) are being checked and detailed. This is also true for the insertable mirror. Parts for the 0F region THz diagnostic were assembled and leak-checked. After integration and alignment of the collection mirror the system will be ready for installation. They will be installed in the X-chamber at the earliest opportunity.

Commissioning activities:
The drive laser can now be synchronized with the accelerator RF. Other EPICS controls for subsystems like the variable attenuator are being resolved.
Other activities:
We prepared for the Industrial Physics Forum tour next week and prepared papers to be presented at the Directed Energy Professionals meeting next month.